the sum of two consecutive odd intergers is not less than -12.
Answer by RAY100(1637)

Let x,,x+2 be consecutive integers
,
not less than,, mean >=
.
(x)+(x+2) >= -12
.
2x +2 >= -12
.
2x >= -14
.
x>= -7,,,,answer
.
check,,,(-7)+(-5)= -12,,ok
.
(-5)+(-3) =-8,,,,ok (>-12),,,,ok
.